# Computational Tradeoffs of Optimization-Based Bound Tightening in ReLU Networks

### 27 Dec 2023 v1

@@ 2.1 Neural Networks
A neural network is a function $f$ defined over a directed graph that maps inputs $x\in\mathbb{R}^{n_{0}}$ to $f(x)\in\mathbb{R}^{n_{L}}$. The directed graph $G=(V,E)$ representing the network decomposes into layers $V=\bigcup_{i\in[L]_{0}}V_{i}$; $V_{0}$ is the input layer and $V_{L}$ as the output layer. Layers $1$ through $L-1$ are the hidden layers. In this work, we consider fully connected feed-forward networks, meaning that all nodes in layer $V_{j}$ are connected to all nodes in layer $V_{j+1}$ and that arcs only connect nodes in adjacent layers. This is mainly to simplify the discussion, as this framework can also be adapted to other feed-forward architectures.
The neural network computes an output in the following way. Each layer $l$ outputs a vector $h^{(l)}$ to the following layer, starting with $h^{(0)}=x$. Then we apply an affine transformation ${W^{(l)}}^{\top}h^{(l)}+b^{(l)}=:a^{(l+1)}$. Finally, the activation function is applied (component-wise) to $a^{(l+1)}$ to produce the output of layer $l+1$, i.e., $h^{(l+1)}=\sigma(a^{(l+1)})$. The last layer often applies an extra transformation, such as $\mbox{softmax}(z)=e^{z_{i}}/\sum_{j=1}^{w}e^{z_{j}}$, to model a probability distribution.
@@ 2.2 Computing Activation Bounds
Note that even weaker bounds can be computed, without the need to solve any optimization problem. This follows from Proposition 1, where upper activation bounds for a node can be computed from those in the previous layer. Lower activation bounds can be computed by reversing the sign of the bounds. If we do this for all the layers iteratively, we will get valid bounds for the entire network.
@@
We remark that in our implementation of Algorithm 1, we parallelized the for loop in line 5; this is possible because problem (2.2) is independent $\forall n\in[n_{m}]$ for a fixed layer $m\in[L]$. Note also that if the MILP solver fails to produce an optimal integer solution, we ask the solver to provide us with the best dual bound available, if there is one, before using the bound from Proposition 1.
@@ 3 Experimental Setup
Third, for each network architecture and each type of activation bounds (strong, weak, naive), we defined ten instances of the verification problem, defined in Section 2.3. For this, we select one network of each architecture with L1 regularization, a pruning threshold $\varepsilon=0.0002$, one random image of each class in MNIST as a reference $x_{0}$ and a radius $\epsilon=0.18$ to perturb the image. The resulting problem was then solved with Gurobi 9.1.2 [20]. A time limit of one hour was used to solve each problem. All runs were performed on Dell PowerEdge C6420 Intel Xeon Gold 6152 @ 2.10 GHz with 22 cores; we allocated 4 CPU cores and 6 GB of RAM per network.
@@ 4.1 Strong Versus Weak Activation Bounds
Our first analysis seeks to determine how different the strong and weak bounds are to each other. This has also been reported in [30], but for the sake of completeness we also include our analysis. For this comparison, we use the relative optimality (RO) gap given by $\frac{|\widehat{LB}-LB|}{|LB|+10^{-10}}$, where $\widehat{LB}$ and $LB$ are the weak and strong lower activation bounds obtained by the LP and MIP, respectively.
In Figure 1, we display a summary of our results. Overall, we can see that the strong and weak bounds are remarkably similar to each other. Though not shown, results are similar for the upper activation bounds. In the worst case, we obtained a weak bound whose value was twice that of the corresponding strong bound (Figure 0(d)). We also observe an expected decay in the strength of the weak bound with respect to the depth of the network. We confirmed this with an analysis of the relative gap layer by layer, which we omit here due to space requirements. However, we remark that we obverse an exponential growth on RO gap by layer.
(a) 1-layer network.
(b) 2-layer network.
(c) 5-layer network.
(d) 10-layer network.
Figure 1: Relative optimality gap on the last layer versus rounding threshold $\varepsilon$ for each architecture and regularization type (L1/L2) and level $\lambda$.
(a) 1-layer network.
(b) 2-layer network.
(c) 5-layer network.
(d) 10-layer network.
Figure 2: Running times of Algorithm 1 for strong and weak bounds in the L1 case.
@@ 4.2 Solving Times
A second comparison we report is the running time (wall clock) on the previous experiments. In Figures 2 and 3, we show the total runtime for the L1 and L2 regularizations, respectively.
(a) 1-layer network.
(b) 2-layer network.
(c) 5-layer network.
(d) 10-layer network.
Figure 3: Running times of Algorithm 1 for strong and weak bounds in the L2 case.
In both regularization types, and for every regularization value $\lambda$, we can see the expected improvement in computation time for the weak bounds. This improvement is much more drastic on deeper networks as shown in Figures 1(c), 1(d), 2(c), 2(d). Also note that the pruning threshold has more impact on the strong bounds than on the weak ones, as shown in Figures 1(b), 1(c), 1(d), 2(b), 2(c), and 2(d). The same phenomenon happens with $\lambda$ in this case.
@@ 4.3 Effectiveness On Verification Problems
(a) Instances that reach optimality.
(b) Instances that reach a gap $\leq 1\%$
@@
In Figure 3(a), we see that strong bounds got the most optimal solutions, followed by weak bounds and naive bounds. This was expected. What is more interesting is that if we take a closer look at the instances that were near optimality (gap lower than $1\%$), as shown in Figure 3(b), we can see that weak bounds have a better performance overall, even surpassing strong bounds in some ranges.

### 30 Jan 2024 v2

@@ 2.1 Neural Networks
A neural network is a function $f$ defined over a directed graph that maps inputs $x\in\mathbb{R}^{n_{0}}$ to $f(x)\in\mathbb{R}^{n_{L}}$. The directed graph $G=(V,E)$ representing the network decomposes into layers $V_{i}$ with $i\in[L]_{0}:=\{0,...,L\}$ such that $V=\bigcup\limits_{i\in[L]_{0}}V_{i}$; $V_{0}$ is the input layer and $V_{L}$ as the output layer. Layers $1$ through $L-1$ are the hidden layers. In this work, we consider fully connected feed-forward networks, meaning that all nodes in layer $V_{j}$ are connected to all nodes in layer $V_{j+1}$ and that arcs only connect nodes in adjacent layers. This is mainly to simplify the discussion, as this framework can also be adapted to other feed-forward architectures.
The neural network computes an output in the following way. Each layer $l$ outputs a vector $h^{(l)}$ to the following layer, starting with $h^{(0)}=x$. Then we apply an affine transformation ${W^{(l)}}^{\top}h^{(l)}+b^{(l)}=:a^{(l+1)}$. Finally, the activation function is applied (component-wise) to $a^{(l+1)}$ to produce the output of layer $l+1$, i.e., $h^{(l+1)}=\sigma(a^{(l+1)})$. The last layer often applies an extra transformation, such as $\mbox{softmax}(z)=e^{z_{i}}/\sum_{j=1}^{w}e^{z_{j}}$, to model a probability distribution. The softmax layer cannot be encoded using MILP, but we can still determine the class by comparing the inputs of softmax since it is a monotonic function.
@@ 2.2 Computing Activation Bounds
Note that even weaker bounds can be computed, without the need to solve any optimization problem. This follows from Proposition 1, where upper activation bounds for a node can be computed from those in the previous layer. Also, by reversing this inequality we have lower activation bounds. If we do this for all the layers iteratively, we will get valid bounds for the entire network.
@@
We remark that in our implementation of Algorithm 1, we parallelized the for loop in line 5; this is possible because problem (2.2) is independent $\forall n\in[n_{m}]$ for a fixed layer $m\in[L]$. Note also that if the MILP solver fails to produce an optimal integer solution, we ask the solver to provide us with the best dual bound available, if there is one, before using the bound from Proposition 1. This could happen in rare cases, for example, if the time limit is reached without a dual bound being reported or if there are numerical issues, among others.
This algorithm could be extended to recurrent neural networks with a finite number of steps, but for the purposes of this paper, we will restrict the scope to feedforward neural networks.
@@ 3 Experimental Setup
Third, for each network architecture and each type of activation bounds (strong, weak, naive), we defined instances of the verification problem from Section 2.3. We didn’t use multiple seeds, but instead we ran ten instances for each configuration to reduce uncertainty. For this, we select one network of each architecture with L1 regularization, a pruning threshold $\varepsilon=0.0002$, one random image of each class in MNIST as a reference $x_{0}$ and a radius $\epsilon=0.18$ to perturb the image. The resulting problem was then solved with Gurobi 9.1.2 [20]. All runs were performed on Dell PowerEdge C6420 Intel Xeon Gold 6152 @ 2.10 GHz with 22 cores; we allocated 4 CPU cores and 6 GB of RAM per network.
Since we would like this study to serve multiple settings, we decided to use a uniform time limit of one hour. For the same reason, we didn’t discount the OBBT running time from the verification problem, since its effect depends on the application in mind and how many times the MILP models are solved. Also, the dataset may be a small one, but we hope that advances and insights from smaller problems can help us tackle much larger ones later, as has happened before in other applications of MIPs.
@@ 4.1 Strong Versus Weak Activation Bounds
Our first analysis seeks to determine how different the strong and weak bounds are to each other. This has also been reported in [30], but for the sake of completeness we also include our analysis. For this comparison, we use the relative bound tightness (RBT):= $\frac{|\widehat{B}-B|}{|B|+10^{-10}}$, where $\widehat{B}$ and $B$ are the weak and strong activation bounds obtained by the LP and MIP, respectively. In Figure 1, we display a summary of our results. Overall, we can see that the strong and weak bounds are remarkably similar to each other. Though not shown, results are similar for the upper activation bounds. In the worst case, we obtained a weak bound whose value was twice that of the corresponding strong bound for net 10x100. We also observe an expected decay in the strength of the weak bound with respect to the depth of the network. We confirmed this with an analysis of the RBT layer by layer, which we omit here due to space requirements. However, we remark that we observe an exponential growth on RBT by layer.
Figure 1: RBT for the lower bound on the last layer versus rounding threshold $\varepsilon$ for each architecture, regularization type (L1/L2) and level $\lambda$.
Figure 2: Running times of Algorithm 1 for strong and weak bounds in the L1 case (top) and the L2 case (bottom).
@@ 4.2 Solving Times
A second comparison we report is the running time (wall clock) on the previous experiments. In Figure 2, we show the total runtime for the L1 and L2 regularization. In both regularization types, and for every regularization value $\lambda$, we can see the expected improvement in computation time for the weak bounds. This improvement is much more drastic on deeper networks for the L1 case and more drastic in shallower networks for the L2 case as shown in Figure 2. Also note that the pruning threshold has more impact on the strong bounds than on the weak ones, as shown in Figure 2 where the curve drops faster for MILPs. The same phenomenon happens with $\lambda$ in this case.
@@ 4.3 Effectiveness On Verification Problems
In Figure 2(b), we see that strong bounds got the most optimal solutions, followed by weak bounds and naive bounds. This was expected. What is more interesting is that if we take a closer look at the instances that were near optimality (gap lower than $1\%$), as shown in Figure 2(a), we can see that weak bounds have a better performance overall, even surpassing strong bounds in some ranges.